OpenStudy (anonymous):

Kate recorded the length of her last six workouts. Workout Length (minutes) 1 45 2 22 3 32 4 54 5 40 6 22 What is the median of the lengths of Kate's workouts? (Round, if necessary.) 35.83 minutes 22 minutes 36 minutes 9 minutes

OpenStudy (anonymous):

The Median is the middle number. Let's put all of the minutes in numerical order 22, 22, 32, 40, 45, 54

OpenStudy (anonymous):

In this case there are 2 middle numbers: 32 and 40 You have to find the average of them (40-32)/2 = 8/2 = 4 32 + 4 = 36 40 - 4 = 36 You answer is 36 :)

OpenStudy (anonymous):

tnx